# Pattern formation and spatiotemporal chaos in relativistic degenerate plasmas

### v1

@@ IV.2 Strong relativistic degeneracy
We consider a more dense regime in which the relativistic degeneracy of electrons becomes more stronger than the regime discussed in Sec. IV.1, i.e., when $R_{0}=25$. This corresponds to the density regime $n_{d0}\sim 9\times 10^{33}~{}\rm{cm}^{-3}$. Since the MI domain expands towards higher values of the modulation wave number $k$ (cf. Fig. 1), the TC and SPC state can coexist at a higher value of $k~{}(<k_{m})\sim 0.21$ than that for the case of $R_{0}=5$.
@@
Figure 6: Initial excitations of master and harmonic solitary patterns for the EM wave field ($|A|$) and the associated density perturbation ($N$) are shown over the spatial domain in the strong relativistic degenerate regime with $R_{0}=25$. It is seen that the pattern selections lead to the excitations of three, seven and sixteen solitary modes corresponding to $k=0.21$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.047$ [subplot (c)] respectively.
@@
Figure 6 shows that the pattern selections with (a) $k=0.21$, (b) $0.11$, and (c) $0.047$ initially lead to the excitations of three, seven, and sixteen solitary modes respectively. While the former (number of excited modes for $k=0.21$) remains the same, the latter two are higher than the modes excited for $R_{0}=5$. It follows that higher the degeneracy effects, larger are the instability domain and the excitations of modes in the pattern selection. As before, the excited peaks of the EM patterns are correlated with the density fluctuations of EAWs. However, the amplitudes of EM waves grow stronger than the density perturbations. It implies that the electron-acoustic wave emission becomes stronger in the strong degeneracy regime. Fig. 7(a) shows that for $k=0.21$, as time progresses, two of the initially excited three patterns propagate incoherently and after a long time they collide and fuse into a new incoherent pattern. The positive values of the largest Lyapunov exponent over the entire domain as in Fig. 8 (a) manifest temporal chaos. However, from Fig. 5 (solid lines) it is seen that the correlation function and the mutual information function do not decay exponentially, and the value of the correlation function lies in $-0.2<C(r)<0.1$. So, the system is in the coexistence of TC and SPC. Although a transfer of energy from initially excited modes to a new incoherent mode occurs, it is not enough to cause STC.
@@
As $k$ is reduced from $k=0.21$ to $k=0.11$, seven solitary patterns are initially excited near $x=\pm L_{x}/8,~{}\pm L_{x}/4,~{}\pm 3L_{x}/8$, and $x=0$ [Fig. 7(b)] due to pattern selection. Two of these coherent patterns, initially peaked near $x=-3L_{x}/8$ and $-L_{x}/4$, collide and fuse into a new incoherent one, and this new pattern collides with the master pattern (initially excited near $x=0$) and gets fused into another new incoherent pattern with a higher amplitude but a reduced wavelength. This newly formed pattern again collides with a pattern that is formed by the fusion of two other patterns that were initially peaked at $x=3L_{x}/8$ and $L_{x}/4$. Finally, seven coherent solitary patterns fuse into a single incoherent distorted pattern with a higher amplitude but a shorter wavelength due to electron-acoustic radiation. In this case, the largest Lyapunov exponent is seen to be positive in the entire spatial domain [Fig. 8(b)]. From Fig. 9 (dashed and dash-dotted lines) it is also clear that both the correlation function and the mutual information function decay exponentially to zero [$C(r)\sim\exp(-r/7.008)$, $I(r)\sim\exp(-r/2.6846)$]. So, the system is in STC in which the spatial redistribution of energy takes place in the process of collision and fusion from initially excited seven modes to a single incoherent mode. It is noted that as the value of the degeneracy parameter $R_{0}$ increases, the contribution from the nonlocal nonlinearity enhances compared to a lower value of $R_{0}$, implying that a few number of initial patterns can also lead to STC by the influence of a strong electron-acoustic wave emission.
@@
In order to exhibit the pattern dynamics of many initially excited modes, we further reduce the modulation wave number to $k=0.047$. It is seen that initially excited sixteen modes [Fig. 6(c)] at different points finally fused into two incoherent patterns after several collisions and fusions among them [Fig. 7(c)]. The pattern initially excited at the boundary $x=-L_{x}/2$ does not propagate any more after $t\approx 250$. However, the eight patterns initially peaked at $x=-nL_{x}/16$ ($n=1,...,7$) and $x=0$ fuse into one new pattern. The remaining seven patterns initially peaked at $x=nL_{x}/16$ ($n=1,...,7$) also collide and eventually fuse into a single new incoherent pattern. Note that in the strong degeneracy regime with reduced values of $k$, the electron-acoustic radiation also becomes stronger (since the stationary states of EM waves gradually disintegrate), which results in faster collisions and fusions among the patterns compared to the moderate degeneracy regime. The largest Lyapunov exponent spectra is seen to be positive [Fig. 8(c)], and the correlation function, and mutual information decay exponentially to zero [$C(r)\sim\exp(-r/3.6443)$, $I(r)\sim\exp(-r/1.5939)$; See the dotted lines of Fig. 9 for $C(r)$ and $I(r)$ and Table 1 for the decay lengths $\xi_{C}$ and $\xi_{I}$]. So, in this case, the system is also in STC. For $R_{0}=25$, the critical wave number ($k_{cs}$) lies in $0.210.11<k_{\rm{cs}}<$ 0.21. So, STC occurs within $0<k<k_{\rm{cs}}$.
IV.3 Ultra-relativistic degeneracy with $R_{0}\sim 85$
@@
From Secs. IV.1 and IV.2, we noted that the state of STC indeed emerges in the nonlinear interaction of EM waves and EAWs in domains of the modulation wave number $(k\ll 1)$ where many unstable modes can be excited initially and strong EAW emission occurs, resulting in faster collisions and fusions among the patterns, as we approach from moderate to strong relativistic degeneracy regimes with increasing values of $R_{0}$. It is also pertinent to examine and verify the existence of STC in the ultra-relativistic degeneracy regime in which $R_{0}\gg 1$. For an illustration purpose, we choose $R_{0}=85$, which corresponds to the density regime $n_{d0}\sim 4\times 10^{35}~{}\rm{cm}^{-3}$. However, it is free to choose any finite value of $R_{0}$ larger than $R_{0}=85$ for which the qualitative features will remain the same. In this ultra-relativistic regime, the modulational instability domain for $k$ is also higher (See the dotted line in Fig. 1) but with a reduced growth rate. This indicates that the system’s energy transfer rate from initially excited many modes to a few modes can be faster compared to the cases of moderate and strong relativistic degeneracy regimes (See Secs. IV.1 and IV.2).
@@
Figure 10: Initial excitations of master and harmonic solitary patterns for the EM wave field ($|A|$) and the associated density perturbation ($N$) are shown over the spatial domain in the ultra-relativistic degenerate regime with $R_{0}=85$. It is seen that the pattern selections lead to the excitations of two, seven and fifteen solitary modes corresponding to $k=0.29$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.047$ [subplot (c)] respectively.
@@
As in the previous cases, we are interested in the domain $0<k<k_{m}$. We first choose a value of $k$, which is close to but below $k_{m}$ and then successively reduce from it for the excitation of many modes due to pattern selection. Since modulational instability domain further expands with a cut-off of the growth rate at higher value of $k$ than that in the case with $R_{0}=25$, we first choose $k=0.29$ to verify the existence of TC and SPC states. From Fig. 10, it is also evident that the EM wave peaks, initially excited at different points, are highly correlated with the density perturbations. Like Fig. 6, peaks of EMWs appear stronger in amplitudes than density perturbations, implying that the EAW emission becomes stronger in higher density regimes of degenerate electrons. For $k=0.29$, initially two solitary patterns are formed [One is the harmonic mode excited near the boundary and another master mode near $x=0$; See Fig. 10(a)]. They collide at time $t\approx 140$ and fuse into a new incoherent pattern [See Fig. 11(a)] with a stronger amplitude but narrower width. The largest Lyapunov exponent in this case is found to be positive in the entire spatial domain of $x$ [See Fig. 12(a)]. However, the correlation function and the mutual information do not decay exponentially to zero [See the solid lines of subplots (a) and (b) of Fig. 13]. Although the energy transfer takes place from initially excited two modes to a single incoherent mode, the collision and fusion are not random to cause STC. Thus, the system is in the coexistence of TC and SPC.
@@
Reducing $k$ from $k=0.29$ to $k=0.11$ gives rise to the excitation of a few more modes than two. We observe that seven patterns are formed initially due to pattern selection [See Fig. 10(b)]. Three patterns initially excited at $x=-L_{x}/8,~{}-L_{x}/4$, and $-3L_{x}/8$ collide and fuse into a new incoherent pattern. Two other pairs of patterns initially excited at $x=0,~{}L_{x}/8$ and $x=L_{x}/4,~{}3L_{x}/8$ collide pairwise and fuse to form two new incoherent patterns. Finally, these three new incoherent patterns again collide at $t\approx 350$ and fuse into a single new incoherent pattern with stronger amplitude but narrower width [See Fig. 11(b)]. The wave energy, which was initially stored in seven unstable modes, is now transferred to a single stable higher harmonic mode with a shorter wavelength [9]. From the analysis of Lyapunov exponent spectra and statistical analysis, it is noted that the largest Lyapunov exponent is positive [See Fig. 12(b)], and the correlation function and mutual information decay exponentially to zero [$C(r)\sim\exp(-r/12.2684)$, $I(r)\sim\exp(-r/3.9002)$; See the dash-dotted lines of subplots (a) and (b) of Fig. 13 and Table 1]. Thus, in the ultra-relativistic degeneracy regime, not only occur the random collisions and fusions among the solitary patterns within a shorter interval, the final energy transfer from seven patterns to a new incoherent pattern also becomes faster than the case with $R_{0}=25$. Evidently, the system is in STC.
@@
In order to illustrate more frequent collisions and fusions among patterns and faster energy transfer at a reduced modulation wave number $k$ compared to the previous case, we consider $k=0.047$. In the latter, fifteen patterns are seen to be excited initially due to pattern selection [See Fig. 10(c)]. From 11(c), we note that the stochastic motions of pattern trains lead to frequent collisions among the neighboring patterns and fuse into new incoherent patterns. Finally, initially excited fifteen patterns fuse into a single incoherent pattern with a higher amplitude but a narrower width at a time shorter than the case with $R_{0}=25$ [See Fig. 7(c)]. From the positive values of the largest Lyapunov exponent over the entire spatial domain [Fig. 12(c)] and the decays of the correlation function and the mutual information (with $\xi_{C}\approx 4.3995$, $\xi_{I}\approx 1.2533$; See Table 1 and the dotted lines of Fig. 13) it is evident that the STC state also emerges in the ultra-relativistic regime. In this case, the critical wave number ($k_{\rm{cs}}$) lies in $0.290.11<k_{\rm{cs}}<$ 0.29.
@@
V Discussion and conclusion

### v2

@@ 4.2 Strong relativistic degeneracy
We consider a more dense regime in which the relativistic degeneracy of electrons becomes more stronger than the regime discussed in Sec. 4.1, i.e., when $R_{0}=25$. This corresponds to the density regime $n_{d0}\sim 9\times 10^{33}~\rm{cm}^{-3}$. Since the MI domain expands with increasing values of $R_{0}$ towards higher values of the modulation wave number $k$ (cf. Fig. 1), the TC and SPC state can coexist at a higher value of $k~(<k_{m})$ than that for the case of $R_{0}=5$. In the case of $R_{0}=25$, the MI domain, where the growth rate tends to become high, is $0<k\lesssim k_{m}\equiv 0.33$ (cf. Fig. 1), and we choose three different values of $k~(<k_{m})$, namely $k=0.21$, $k=0.11$, and $k=0.047$ for the excitation of increasing number of wave modes ($k^{-1}]$) at different wavelengths due to the pattern selection.
@@
Figure 6: Initial excitations of master and harmonic solitary patterns for the EM wave field ($|A|$) and the associated density perturbation (NN) are shown over the spatial domain in the strong relativistic degenerate regime with $R_{0}=25$. It is seen that the pattern selections lead to the excitations of three, seven and sixteen solitary modes corresponding to $k=0.21$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.047$ [subplot (c)] respectively. The other fixed parameter values are $A_{0}=0.8$ and $b=0.5$.
@@
Figure 6 shows that the pattern selections with (a) $k=0.21$, (b) $0.11$, and (c) $0.047$ initially lead to the excitations of three, seven, and sixteen solitary modes respectively. While the former (number of excited modes for $k=0.21$) remains the same, the latter two are higher than the modes excited for $R_{0}=5$. It follows that higher the degeneracy effects, larger are the instability domain and the excitations of modes in the pattern selection. As before, the excited peaks of the EM patterns are correlated with the density fluctuations of EAWs. However, the amplitudes of EM waves grow stronger than the density perturbations. It implies that the electron-acoustic wave emission becomes stronger in the strong degeneracy regime. Fig. 7(a) shows that for $k=0.21$, as time progresses, two of the initially excited three patterns propagate incoherently and after a long time they collide and fuse into a new incoherent pattern. The positive values of the largest Lyapunov exponent over the entire domain as in Fig. 8 (a) manifest temporal chaos. However, from Fig. 5 (solid lines) it is seen that the correlation function and the mutual information function do not decay exponentially, and the value of the correlation function lies in 0.2<C(r)<0.1-0.2<C(r)<0.1. So, the system is in the coexistence of TC and SPC. Although a transfer of energy from initially excited modes to a new incoherent mode occurs, it is not enough to cause STC.
@@
As $k$ is reduced from $k=0.21$ to $k=0.11$, seven solitary patterns are initially excited near $x=\pm L_{x}/8,~\pm L_{x}/4,~\pm 3L_{x}/8$, and $x=0$ [Fig. 7(b)] due to pattern selection. Two of these coherent patterns, initially peaked near $x=-3L_{x}/8$ and $L_{x}/4$, collide and fuse into a new incoherent one, and this new pattern collides with the master pattern (initially excited near $x=0$) and gets fused into another new incoherent pattern with a higher amplitude but a reduced wavelength. This newly formed pattern again collides with a pattern that is formed by the fusion of two other patterns that were initially peaked at $x=3L_{x}/8$ and $L_{x}/4$. Finally, seven coherent solitary patterns fuse into a single incoherent distorted pattern with a higher amplitude but a shorter wavelength due to electron-acoustic radiation. In this case, the largest Lyapunov exponent is seen to be positive in the entire spatial domain [Fig. 8(b)]. From Fig. 9 (dashed and dash-dotted lines) it is also clear that both the correlation function and the mutual information function decay exponentially to zero $C(r)\sim\exp(-r/7.008)$, $I(r)\sim\exp(-r/2.6846)$]. So, the system is in STC in which the spatial redistribution of energy takes place in the process of collision and fusion from initially excited seven modes to a single incoherent mode. It is noted that as the value of the degeneracy parameter $R_{0}$ increases, the contribution from the nonlocal nonlinearity enhances compared to a lower value of $R_{0}$, implying that a few number of initial patterns can also lead to STC by the influence of a strong electron-acoustic wave emission.
@@
In order to exhibit the pattern dynamics of many initially excited modes, we further reduce the modulation wave number to $k=0.047$. It is seen that initially excited sixteen modes [Fig. 6(c)] at different points finally fused into two incoherent patterns after several collisions and fusions among them [Fig. 7(c)]. The pattern initially excited at the boundary $x=-L_{x}/2$ does not propagate any more after $t\approx 250$. However, the eight patterns initially peaked at $x=-nL_{x}/16$ (n=1,,7n=1,...,7) and $x=0$ fuse into one new pattern. The remaining seven patterns initially peaked at $x=nL_{x}/16$ (n=1,,7n=1,...,7) also collide and eventually fuse into a single new incoherent pattern. Note that in the strong degeneracy regime with reduced values of $k$, the electron-acoustic radiation also becomes stronger (since the stationary states of EM waves gradually disintegrate), which results in faster collisions and fusions among the patterns compared to the moderate degeneracy regime. The largest Lyapunov exponent spectra is seen to be positive [Fig. 8(c)], and the correlation function, and mutual information decay exponentially to zero $C(r)\sim\exp(-r/3.6443)$, $I(r)\sim\exp(-r/1.5939)$; See the dotted lines of Fig. 9 for $C(r)$ and $I(r)$ and Table 2 for the decay lengths $\xi_{C}$ and $\xi_{I}$]. So, in this case, the system is also in STC. For $R_{0}=25$, the critical wave number ($k_{cs}$) lies in $0.11<k_{\rm{cs}}<0.21$. So, STC occurs within $0<k<k_{\rm{cs}}$.
4.3 Ultra-relativistic degeneracy with $R_{0}\sim 85$
@@
From Secs. 4.1 and 4.2, we noted that the state of STC indeed emerges in the nonlinear interaction of EM waves and EAWs in domains of the modulation wave number $(k\ll 1)$ where many unstable modes can be excited initially and strong EAW emission occurs, resulting in faster collisions and fusions among the patterns, as we approach from moderate to strong relativistic degeneracy regimes with increasing values of $R_{0}$. It is also pertinent to examine and verify the existence of STC in the ultra-relativistic degeneracy regime in which $R_{0}\gg 1$. For an illustration purpose, we choose $R_{0}=85$, which corresponds to the density regime $n_{d0}\sim 4\times 10^{35}~\rm{cm}^{-3}$. However, it is free to choose any finite value of $R_{0}$ larger than $R_{0}=85$ for which the qualitative features will remain the same. In this ultra-relativistic regime, the modulational instability domain for $k$, where the growth rate tends to become high is also higher, i.e., $0<k<k_{m}\equiv 0.62$ (See the dotted line in Fig. 1) but with a reduced growth rate. This indicates that the system’s energy transfer rate from initially excited many modes to a few modes can be faster compared to the cases of moderate and strong relativistic degeneracy regimes (See Secs. 4.1 and 4.2).
@@
Figure 10: Initial excitations of master and harmonic solitary patterns for the EM wave field ($|A|$) and the associated density perturbation (NN) are shown over the spatial domain in the ultra-relativistic degenerate regime with $R_{0}=85$. It is seen that the pattern selections lead to the excitations of two, seven and fifteen solitary modes corresponding to $k=0.29$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.047$ [subplot (c)] respectively. The other fixed parameter values are $A_{0}=0.54$ and $b=0.4$.
@@
As in the previous cases, we are interested in the domain $0<k<k_{m}\equiv 0.44$. We first choose a value of $k$, which is close to but below $k_{m}$ and then successively reduce from it for the excitation of many modes due to pattern selection. Since modulational instability domain further expands with a cut-off of the growth rate at higher value of $k$ than that in the case with $R_{0}=25$, we first choose $k=0.29$ to verify the existence of TC and SPC states, and then successively reduce it to consider the values, $k=0.11$ and $k=0.047$. From Fig. 10, it is also evident that the EM wave peaks, initially excited at different points, are highly correlated with the density perturbations. Like Fig. 6, peaks of EMWs appear stronger in amplitudes than density perturbations, implying that the EAW emission becomes stronger in higher density regimes of degenerate electrons. For $k=0.29$, initially two solitary patterns are formed [One is the harmonic mode excited near the boundary and another master mode near $x=0$; See Fig. 10(a)]. They collide at time $t\approx 140$ and fuse into a new incoherent pattern [See Fig. 11(a)] with a stronger amplitude but narrower width. The largest Lyapunov exponent in this case is found to be positive in the entire spatial domain of $x$ [See Fig. 12(a)]. However, the correlation function and the mutual information do not decay exponentially to zero [See the solid lines of subplots (a) and (b) of Fig. 13]. Although the energy transfer takes place from initially excited two modes to a single incoherent mode, the collision and fusion are not random to cause STC. Thus, the system is in the coexistence of TC and SPC.
@@
Reducing $k$ from $k=0.29$ to $k=0.11$ gives rise to the excitation of a few more modes than two. We observe that seven patterns are formed initially due to pattern selection [See Fig. 10(b)]. Three patterns initially excited at $x=-L_{x}/8,~-L_{x}/4$, and $3L_{x}/8$ collide and fuse into a new incoherent pattern. Two other pairs of patterns initially excited at $x=0,~L_{x}/8$ and $x=L_{x}/4,~3L_{x}/8$ collide pairwise and fuse to form two new incoherent patterns. Finally, these three new incoherent patterns again collide at $t\approx 350$ and fuse into a single new incoherent pattern with stronger amplitude but narrower width [See Fig. 11(b)]. The wave energy, which was initially stored in seven unstable modes, is now transferred to a single stable higher harmonic mode with a shorter wavelength [he2002harmonic]. From the analysis of Lyapunov exponent spectra and statistical analysis, it is noted that the largest Lyapunov exponent is positive [See Fig. 12(b)], and the correlation function and mutual information decay exponentially to zero $C(r)\sim\exp(-r/12.2684)$, $I(r)\sim\exp(-r/3.9002)$; See the dash-dotted lines of subplots (a) and (b) of Fig. 13 and Table 2]. Thus, in the ultra-relativistic degeneracy regime, not only occur the random collisions and fusions among the solitary patterns within a shorter interval, the final energy transfer from seven patterns to a new incoherent pattern also becomes faster than the case with $R_{0}=25$. Evidently, the system is in STC.
@@
In order to illustrate more frequent collisions and fusions among patterns and faster energy transfer at a reduced modulation wave number $k$ compared to the previous case, we consider $k=0.047$. In the latter, fifteen patterns are seen to be excited initially due to pattern selection [See Fig. 10(c)]. From 11(c), we note that the stochastic motions of pattern trains lead to frequent collisions among the neighboring patterns and fuse into new incoherent patterns. Finally, initially excited fifteen patterns fuse into a single incoherent pattern with a higher amplitude but a narrower width at a time shorter than the case with $R_{0}=25$ [See Fig. 7(c)]. From the positive values of the largest Lyapunov exponent over the entire spatial domain [Fig. 12(c)] and the decays of the correlation function and the mutual information (with $\xi_{C}\approx 4.3995$, $\xi_{I}\approx 1.2533$; See Table 2 and the dotted lines of Fig. 13) it is evident that the STC state also emerges in the ultra-relativistic regime. In this case, the critical wave number ($k_{\rm{cs}}$) lies in $0.11<k_{\rm{cs}}<0.29$.
Relativistic Degeneracy ($R_{0}$) MI domain ($\Gamma$ tends to become high) Large-scale excitation ($0<k\lesssim k_{m}$) Pattern selection $\left(M=\left[k^{-1}\right]\right)$; Excitation of increasing number of EM solitary waves at different $k<k_{m}$ Appearance of different states Critical wave number ($k_{cs}$) at which transition from one state to another occurs
$k=0.15$ TC & SPC
$k=0.11$ STC
Moderate: $R_{0}=5$, i.e., $n_{\rm{d0}}\sim 8\times 10^{31}~\rm{cm}^{-3}$ $0<k\lesssim k_{m}\equiv 0.18$ $k=0.043$ STC $k_{cs}\in\left(0.11,0.15\right)$
$k=0.21$ TC & SPC
$k=0.11$ STC
Strong: $R_{0}=25$, i.e., $n_{\rm{d0}}\sim 9\times 10^{33}~\rm{cm}^{-3}$ $0<k\lesssim k_{m}\equiv 0.33$ $k=0.047$ STC $k_{cs}\in\left(0.11,0.21\right)$
$k=0.29$ TC & SPC
$k=0.11$ STC
Ultra-relativistic: $R_{0}=85$, i.e., $n_{\rm{d0}}\sim 4\times 10^{35}~\rm{cm}^{-3}$ $0<k\lesssim k_{m}\equiv 0.44$ $k=0.047$ STC $k_{cs}\in\left(0.11,0.29\right)$
Table 1: A comparative summary showing the emergence of different states of the system in three different regimes of the relativistic degeneracy ($R_{0}$).
@@
To compare the results in Secs. 4.14.3 across three degeneracy regimes, we present a summary in Table 1. As the degeneracy regime changes from moderate to strong to ultra-relativistic with increasing $R_{0}$, the intervals admitting the critical value $k_{\rm{cs}}$ expand. As a result, the domains of $k$ for both the coexistence of TC and SPC states, as well as the presence of STC states, also increase. Thus, the STC state appears in a wider domain of $k$ under ultra-relativistic degeneracy than under moderate or strong degeneracy. Since the MI domain grows with increasing $R_{0}$, shifting from moderate to ultra-relativistic regimes reduces the possibility of stable EM solitons but enhances the emergence of STC states.
5 Discussion and conclusion